OpenStudy (anonymous):

A sample of zinc with a mass of 130.8g contains 6.02 x 10^23 atoms. Calculate the mass of a single zinc atom.

OpenStudy (anonymous):

\[\frac{ Zn }{ 1 }\times \frac{ grams of one Zn }{ Zn atoms }\] Without numbers, can you figure it out?

OpenStudy (anonymous):

gramsofonezn =grams of one zinc Znatoms= zinc atoms.

OpenStudy (anonymous):

@Shalante can i try it out and then ask you if it's right? D:

OpenStudy (anonymous):

Sure! I answered your question by the way.

OpenStudy (anonymous):

Sorry new to this and okay thanks!

OpenStudy (anonymous):

Oh wait. I thought you were on a different board. I got confused my bad.

OpenStudy (anonymous):

I gtg. Be back in 90minutes 1Zn x (130.8g/6.023 x 10^23 Zn atoms.

OpenStudy (anonymous):

okay thanks

OpenStudy (jhannybean):

\[\sf 130.8~g~Zn ~\times ~\frac{1~mol~ Zn}{65.38 ~g~Zn}~\times \frac{6.02~\times~10^{23}~Zn~atoms}{1~mol~Zn} \]\[\sf = \frac{(130.8 \cdot 6.02~\times~10^{23})}{65.38}~Zn ~atoms \] :)

OpenStudy (anonymous):

@Jhannybean wait where did you get 65.38 from

OpenStudy (jhannybean):

Molar mass of zinc

OpenStudy (anonymous):

ahhh im confused

OpenStudy (anonymous):

Hmm, did I make a mistake? I was helping 4 people at the same time.

OpenStudy (anonymous):

I dont think we need molar mass since the grams would cancel out and we need to know the mass of a single Zinc atoms.

OpenStudy (anonymous):

@Jhannybean Your answer is in Zn atoms. Isnt it supposed to be the gram of one atom?

OpenStudy (anonymous):

Maybe it is 130.8g Zn/65.38g Zn=2.00 mol Zn \[\frac{ 2.00 mol Zn }{ 1 }\times \frac{ 6.022\times10^{23} Zn atoms }{ 1mol Zn }=1.2044\times10^{24}Zn atoms\]=1.2044 x 10^24 Zn atoms Divide 130.8 grams by that big number.

OpenStudy (anonymous):

Not sure why the question tells us 130.8 g of Zn contains 6.022 x 10^23 atoms. (Is it trying to trick us or is it an exception?)

OpenStudy (anonymous):

I think my first formula is right.

OpenStudy (anonymous):

If not, you can blame your teacher because of the bad wording.